OpenStudy (anonymous):

Write the equation of the line that passes through (2, 2) and (6, 3) in standard form. 4x + 4y = 6 x + 4y = 12 4x + 4y = –12 x – 4y = –6

OpenStudy (anonymous):

lets find the slope(m)... m = (y2 - y1) / (x2 - x1) set 1 (2,2) set 2 (6,3) m = (3 - 2) / (6 - 2) m = 1/4 (this is the slope) Now we put it in this form : y = mx + b you can use either of your points to represent x and y...I will use (2,2) y = mx + b 2 = 1/4(2) + b 2 = 1/2 + b 2 - 1/2 = b 2/2 - 1/2 = b 1/2 = b your equation is y = 1/4x + 1/2 now we will put it in standard form :Ax + By = C -1/4x + y = 1/2 (multiply the equation by -4) x - 4y = - 2 <-- equation in standard form